OpenStudy (anonymous):

What is the percent composition of water, H2O? no idea how to do this at all, please help. a.)11.2% H, 88.8% O b.)1.01% H, 16.0% O c.)5.59% H, 94.4% O d.)20.2% H, 79.8% O e.)66.7% H, 33.3% O

OpenStudy (cuanchi):

H%= 2 x 100 / 18 = ???? O%= 16 x 100 /18 = ???

OpenStudy (frank0520):

A

rebeccaxhawaii (rebeccaxhawaii):

\(\Huge\boxed{\boxed{\boxed{\huge\mathbb{WELCOME~TO~OPENSTUDY} }}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with your studies! \(\Huge\heartsuit\)